a particle is projected at an angle of 30 to the horizontal and 2 sec later is moving in direction tan -1 (1/4) to the horizontal for the first time . find its initial speed

Grade:11

1 Answers

Sawan raj
4 years ago
At the point where the angle is tan-1/4 the component of velocity in x direction will be ucos 30 and velocity in y direction will be usin30 - gt so it is eqaul ro usin30- 20
            Tan-1/4 =usin30-20/ucos30
         1/4 =usin30-20/ucos 30
         U= 160/4-1.732